1. An argon-ion laser emits a blue-green beam of light with a wavelength of 488nm in a vacuum.
Compute the energy in joules for the two energy states for the atomic transition which produces
this light?
1
Expert's answer
2020-06-08T10:14:20-0400

"v = \frac{c}{\lambda}"


"c=3\times10^{8}ms^{-1}"


"\lambda = 488 \times 10 ^ {-9}=4.88 \times 10^{-7}m"


"Energy = \frac{3\times10^{8}}{4.88 \times 10^{-7}} = 6.14\times 10 ^{14}"